OpenStudy (anonymous):

Anybody need help with fractions or turning fractions into something like:decimals,improper fractions come straight to me and i'll see what i can do!!!

OpenStudy (anonymous):

Which statement is true? Converting an improper fraction to a decimal shows if it is rational. Repeating decimals are never rational. A zero numerator over an integer denominator is never rational. A negative integer is always rational.

OpenStudy (anonymous):

which one is it??

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (solomonzelman):

@happyface7 A, as I said before.

OpenStudy (solomonzelman):

It is D @izzy_swaggman. B is false

OpenStudy (solomonzelman):

Sorry for this mistake!

OpenStudy (anonymous):

how is that true @SolomonZelman

OpenStudy (solomonzelman):

Well you can eliminate Converting an improper fraction to a decimal shows if it is rational. Obviously, lets say you have 4/73. Repeating decimals are never rational. repeating decimals are always rational. A zero numerator over an integer denominator is never rational. 0/3=0, by definition zero is a rational number https://www.google.com/search?q=rational+numbers&source=lnms&tbm=isch&sa=X&ei=g4CCUrGhDNGtkAf9w4GADg&sqi=2&ved=0CAcQ_AUoAQ&biw=1920&bih=955#facrc=_&imgdii=_&imgrc=2E3Xm11PasI5WM%3A%3BS24H-C_xfjYcUM%3Bhttp%253A%252F%252Fscimathmn.org%252Fstemtc%252Fsites%252Fdefault%252Ffiles%252Fimages%252Fframeworks%252Fmath%252F8.1.1A%252Fimage131.jpg%3Bhttp%253A%252F%252Fscimathmn.org%252Fstemtc%252Fframeworks%252F811a-rational-irrational-real-numbers%3B866%3B386 A negative integer is always rational. look at the third word...

OpenStudy (solomonzelman):

Cool!